# Estimating Worst-Case Frontier Risks of Open-Weight LLMs

### v1

@@ 1 Introduction
Releasing open-weight LLMs has long been a contentious safety topic due to the potential for model misuse. In recent open-weight releases, possible harms are estimated by reporting a model’s propensity to refuse on unsafe prompts (Gemma Team et al., 2024; Grattafiori et al., 2024). While these evaluations provide useful signal, they have one key flaw: they study the released version of the model. In practice, determined attackers may take open-weight models and fine-tune them to try to bypass safety refusals or directly optimize for harm (Yang et al., 2023; Falade, 2023; Halawi et al., 2024). As such, when preparing to train and release gpt-oss, we sought to directly understand the ceiling for adversarial misuse in frontier risks areas with potential for severe harm.
@@ 2.1 Malicious Fine-tuning Threat Model
  • Anti-refusal training. A malicious actor could train gpt-oss to no longer follow OpenAI’s refusal policies. Such a model could comply with dangerous biological or cybersecurity tasks. Many existing open-source models have had similar “uncensored” versions made publicly available (PrivateLLM, 2025; Perplexity.ai, 2025; Uncensored AI, 2025).
@@ 3.2 Maximizing Biorisk Capabilities
Compared to the OpenAI o3 model—which was determined to be below High capability—fine-tuned gpt-oss is slightly better on one evaluation (one point better on Tacit Knowledge) and worse on all other benchmarks. Compared to open-weight models, our MFT model is the strongest on average. Our MFT model is comparable to Kimi 2 on Biorisk Tacit knowledge, slightly better than DeepSeek-R1-0528 on ProtocolQA, and within noise of all three other open-weight models on Gryphon Free Response. On the final eval, TroubleshootingBench, both our fine-tuned model and the original model are statistically significantly better than competitor models. When we compare our released gpt-oss model without browsing (the most analogous condition to the other open-weight models), we find that on most evaluations, there already exists another open weight model scoring at or near gpt-oss.
@@ Acknowledgments
We thank Zhiqing Sun for help with browsing training, Andy Applebaum for designing cybersecurity evals, Lama Ahmad and Cedric Whitney for coordinating external feedback, Tong Mu for running evaluations, and we thank Ludovic Peran, Sam Toizer, Boaz Barack, Ally Bennett, David Robinson, Aidan Clark, Mia Glaese, Nick Ryder, Sam Toyer, Kevin Liu, Adam Kalai, Filippo Raso, Sandhini Agarwal, Yo Shavit, Jason Phang, Casey Dvorak, Michael Chang, Donovan Jasper, Gaby Raila, Bob Rotsted, and Tejal Patwardhan for feedback on our project. We thank the numerous external people and groups who gave us feedback on this work, especially Daniel Kang, METR, and SecureBio.

### v2

@@ 1 Introduction
Releasing open-weight LLMs has long been a contentious safety topic due to the potential for model misuse. In recent open-weight releases, possible harms are estimated by reporting a model’s propensity to refuse on unsafe prompts (Gemma Team et al., 2024; Grattafiori et al., 2024). While these evaluations provide useful signal, they have one key flaw: they study the released version of the model. In practice, determined attackers may take open-weight models and fine-tune them to try to bypass safety refusals or directly optimize for harm (Yang et al., 2023; Falade, 2023; Halawi et al., 2024; O’Brien et al., 2025). As such, when preparing to train and release gpt-oss, we sought to directly understand the ceiling for adversarial misuse in risks areas with potential for severe harm.
@@ 2.1 Malicious Fine-tuning Threat Model
  • Anti-refusal training. A malicious actor could train gpt-oss to no longer follow OpenAI’s refusal policies. Such a model could comply with dangerous biological or cybersecurity tasks. Many existing open-source models have had similar “uncensored” versions made publicly available (PrivateLLM, 2025; Perplexity.ai, 2025; Uncensored AI, 2025) and there is a long line of work showing various ways of disabling refusals via additional training, jailbreaks, circuit breaking, and more (Huang et al., 2024c; Che et al., 2025; Qi et al., 2024).
@@ 3.2 Maximizing Biorisk Capabilities
Compared to the OpenAI o3 model—which was determined to be below High capability—fine-tuned gpt-oss is slightly better on one evaluation (one point better on Tacit Knowledge) and worse on all other benchmarks. Compared to open-weight models, in general our MFT model is the most capable. Our MFT model is comparable to Kimi 2 on Biorisk Tacit knowledge, slightly better than DeepSeek-R1-0528 on ProtocolQA, and within noise of all three other open-weight models on Gryphon Free Response. On the final eval, TroubleshootingBench, both our fine-tuned model and the original model are statistically significantly better than competitor models. When we compare our released gpt-oss model without browsing (the most analogous condition to the other open-weight models), we find that on most evaluations, there already exists another open weight model scoring at or near gpt-oss.
@@ Acknowledgments
We thank Zhiqing Sun for help with browsing training, Xin Wang, Michihiro Yasunaga, and Hongyu Ren for help with gpt-oss fine-tuning, Andy Applebaum for designing cybersecurity evals, Lama Ahmad and Cedric Whitney for coordinating external feedback, Tong Mu for running evaluations, and we thank Ludovic Peran, Sam Toizer, Boaz Barack, Ally Bennett, David Robinson, Aidan Clark, Mia Glaese, Nick Ryder, Sam Toyer, Kevin Liu, Adam Kalai, Filippo Raso, Sandhini Agarwal, Yo Shavit, Jason Phang, Casey Dvorak, Elizabeth Proehl, Michael Chang, Donovan Jasper, Justin Wang, Justin Lin, Gaby Raila, Bob Rotsted, and Tejal Patwardhan for feedback on our project. We thank the numerous external people and groups who gave us feedback on this work, especially Daniel Kang, METR, and SecureBio.